the sum o two numbers is 24. the sum of the greater.number and twice the smaller is 18 find the greatest number
Answer by stanbon(75887)

the sum o two numbers is 24. the sum of the greater.number and twice the smaller is 18 find the greatest number
----------------------
x + y = 24
x + 2y = 18
----
Subtract and solve for "y"::
y = -6
------
x + y = 24
x -6 = 24
x = 30
-----------------
Cheers,
Stan H.
-------------